# OpenScan: A Benchmark for Generalized Open-Vocabulary 3D Scene Understanding

### v2

@@ Abstract
1 Introduction
Open-vocabulary 3D scene understanding (OV-3D) involves recognizing object classes that are not included in the training set. It is important to applications like autonomous driving [4] and robotics [42]. Recently, vision-language models (VLMs), e.g., CLIP [27], have achieved significant progress by leveraging large-scale image-text datasets with semantically rich captions. The impressive capability of VLMs in capturing rich contexts between images and texts has inspired further exploration in open-vocabulary tasks in both 2D [15, 43] and 3D [34, 25] domains.
For AI systems, the capability to comprehend diverse linguistic aspects of object-related attributes and their association with corresponding objects is as important as the identification of the objects themselves. Consequently, the field of open-vocabulary 3D scene understanding should ideally extend beyond specific object classes to encompass complex object-related attributes, such as affordance and material, articulated through natural languages. However, the generalization ability of existing OV-3D methods [25, 34, 36, 41, 24] to object-related attributes has not been thoroughly and systematically explored. Besides, evaluating the ability of an OV-3D model to recognize specific object attributes is difficult due to the shortage of large-scale OV-3D attribute benchmarks. Existing OV-3D benchmarks, such as ScanNet [8] and ScanNet200 [30], primarily focus on object classes, and do not include annotations of object attributes for evaluating the generalized ability of OV-3D methods.
@@ 1 Introduction
| ScanNet | ScanNet200 | ScanNet++ | OpenScan | |
| Scan Source | ScanNet | ScanNet | ScanNet++ | ScanNet |
| Object Annotation | Class | Class | Class | Attribute |
Existing 3D scene understanding benchmarks, such as ScanNet [8], ScanNet200 [30], and ScanNet++ [40], only provide annotations for object classes. To address this limitation of existing benchmarks, we construct a new benchmark, named OpenScan, for the GOV-3D task. OpenScan is constructed based on the ScanNet200 [30] benchmark. It expands the single category of object classes in ScanNet200 into eight linguistic aspects of object-related attributes, including affordance, property, type, manner, synonym, requirement, element, and material. This allows each object to be associated with some generalized knowledge beyond object classes. With our OpenScan benchmark, it becomes possible to comprehensively evaluate existing OV-3D models from various aspects, enabling a quantitative assessment of their generalization capabilities in understanding abstract object attributes.
@@
2 Related Work
Open-Vocabulary 2D Understanding Benchmarks. Open-vocabulary 2D understanding refers to the task of detecting or segmenting novel object classes that are not present in the training dataset. For the object detection task, COCO [20] and LVIS [16] are two widely used datasets. For the image segmentation task, popular datasets include COCO [20], ADE20k [44], PASCAL-VOC [12], and Cityscapes [7]. However, these benchmarks primarily evaluate the model’s open-vocabulary ability but do not explicitly assess its capability to recognize specific object characteristics. PACO [28] introduces a 2D segmentation benchmark that focuses on parts and attributes of common objects. Inspired by PACO [28], FG-OVD [3] presents a challenging task and benchmark for fine-grained open-vocabulary object detection to evaluate the ability of open-vocabulary detectors to discern extrinsic object properties. Similarly, OVDEval [38] introduces an open-vocabulary detection benchmark to evaluate the performance on linguistic aspects using complex language prompts. Our work is different from them [28, 3, 38] since we focus on the understanding of object attributes on 3D data, which poses greater challenges compared to the understanding in 2D images due to the limited annotations in 3D benchmarks.
@@ 2 Related Work
Open-Vocabulary 3D Scene Understanding. The study of open-vocabulary 3D scene understanding has been relatively limited compared to open-vocabulary 2D understanding. This is primarily due to the complexity and difficulty in obtaining 3D datasets. OpenMask3D [34] introduces the open-vocabulary 3D instance segmentation task. It proposes the first approach for the task in a zero-shot setting. OpenScene [25] also proposes a zero-shot method for open-vocabulary 3D scene understanding. Beyond the object class, it is able to utilize arbitrary text queries for semantic segmentation. Previous methods have mainly focused on object context for 3D scene understanding. PLA [10] and RegionPLC [37] extend the context to a more coarse-to-fine semantic representation to provide a more comprehensive supervision. Recently, OpenIns3D [17], Open3DIS [24], and SAI3D [41] utilize powerful 2D segmentation models to generate 2D instances and then merge them into 3D instances. Instead of utilizing accurate 2D masks from 2D segmentation models, MaskClustering [36] leverages clustering algorithms to perform zero-shot 3D segmentation. Recently, UniSeg3D [35] proposes a unified framework for 3D scene understanding. However, these methods only provide qualitative results for object attributes and lack a thorough evaluation of performance beyond object classes. This motivates us to conduct a quantitative evaluation that encompasses a wider range of object attributes.
@@
Open-Vocabulary 3D Scene Understanding Benchmark. Existing open-vocabulary 3D scene understanding benchmarks, e.g., ScanNet [8], ScanNet200 [30], S3DIS [1], and Matterport3D [5], utilize RGB-D cameras, while ARKitScenes [2], Replica [33], and ScanNet++ [40] leverage high-resolution laser scanners to capture high-fidelity 3D data for 3D reconstructions. Recently, SceneFun3D [9] provides a large-scale 3D dataset with functionality and affordance interaction annotations in 3D scenes. However, our OpenScan differs from SceneFun3D by considering a broader range of attributes. Specifically, while SceneFun3D focuses on function or affordance understanding, our OpenScan covers eight linguistic aspects, with affordance representing just one of them. Besides, while SceneFun3D focuses on element-level human-scene interaction (e.g., “door handle”), our OpenScan is designed for object-level scene understanding (e.g., “door”).
3 Task Setting and Benchmark
3.1 Task Formulation
@@
OV-3D. Let represent 3D scenes with points and denote a vocabulary set composed of text sentences, each describing the object class that we aim to detect. An OV-3D model, , generates predictions with high confidence scores, . Predictions are then compared with the GT label for evaluation.
@@
GOV-3D. The existing 3D scene understanding benchmark, denoted as , comprises a collection of object-label pairs. Each pair consists of an object represented as a point cloud and its corresponding class label . The benchmark is composed of multiple 3D scenes with points. Building upon this, GOV-3D extends the class label to object attribute . Accordingly, the attribute set for 3D scenes is defined as , composed of text sentences, each describing a specific attribute that we aim to detect. A GOV-3D model, , produces predictions with high confidence scores, . The evaluation of the GOV-3D task involves comparing the predictions with the GT label .
Metrics. We employ commonly used OV-3D metrics to evaluate our GOV-3D task. For semantic segmentation, we follow [25, 10, 37] to apply mean IoU (mIoU) and mean accuracy (mAcc). For instance segmentation, we follow [34, 41, 36, 24] to apply average prevision (AP) at IoU scores of 25% (AP25), 50% (AP50), and the mean of AP from 50% to 95% at 5% steps.
3.2 Benchmark Description
@@
The OpenScan benchmark is constructed based on the ScanNet200 [30] benchmark, which consists of 200 object classes with more than 1,500 3D scans. Since the ScanNet200 benchmark contains only an object-level class annotation for each object, it is not suitable for our GOV-3D task. To perform the GOV-3D task, we construct the OpenScan benchmark by leveraging the object annotations of the ScanNet benchmark. Our OpenScan provides attribute annotations for each object, expanding the single category of object classes in ScanNet200 into eight linguistic aspects of object-related attributes, including affordance, property, type, manner, synonym, requirement, element, and material. Figure 2 shows an example from our OpenScan benchmark. The target objects in our OpenScan are annotated with eight linguistic aspects of object attributes. The explanations of these eight object attributes are as follows:
@@
3.3 Benchmark Annotation
Object-Attribute Association with Knowledge Graph. We associate each object with various attributes using knowledge graphs, as illustrated in Figure 3. Let denotes the existing 3D scene understanding benchmark, e.g., ScanNet200 [30] in our implementation, where is a target object, is the corresponding class label, and denotes the number of target object and label pairs. The benchmark is composed of multiple 3D scenes with points. Let denotes the knowledge graph, where is the node set and is the edge set. The nodes are natural language words and phrases, and the edges are relation knowledge connecting them. Each edge is directional, and can be represented as a tuple , where are the names of the head node and the tail node, is the relation, and is the importance weight of this relation. We extract the relation knowledge from the popular and high-quality NLP knowledge base ConceptNet [32]. An example of relation knowledge from it is:
We query a set of relation knowledge that relates to an input object class from the knowledge graph . Formally, for each edge within it, the head node name is the same as the input object class, i.e., . The query process can be formulated as:
These object-attribute pairs constitute the basic data annotation of our OpenScan benchmark, which is useful in the GOV-3D task. Finally, each 3D object is assigned a relation knowledge through annotations, serving as commonsense knowledge as:
Manual Annotation. For the visual attribute that cannot be inferred without human perception, we manually annotate the attribute for each 3D target object in our benchmark following a rigorous protocol. We create a web interface for the annotators to select the corresponding visual attribute of each target object. Specifically, for each scene, annotators are presented with the 3D point cloud and the corresponding 2D image frame of the target object. Taking the material attribute as an example, annotators are tasked with identifying the primary material composition of the target object. Any 3D object with an ambiguous appearance is carefully identified through different camera views of the scene and the corresponding image frames around the object. Finally, each 3D object is assigned with a relation and a visual attribute like material through annotations, serving as visual appearance in:
Attribute Classification. To better organize our benchmark, we manually classify each object attribute into eight linguistic aspects according to the relation and attribute . This process involves considering the nature of the relation and attribute , and how they align with each linguistic aspect. Subsequently, each attribute is assigned to a linguistic aspect.
Attribute Verification. After the initial attribute classification, we manually verify each 3D object with its corresponding attribute and linguistic aspect. If a 3D object contains multiple attributes within a single linguistic aspect, we manually select the most representative one to avoid inconsistency during evaluation. We also filter out similar attributes (e.g., “store things” and “store somethings”), preserving only one attribute to ensure consistency.
Query Generation. A practical query in our GOV-3D task should incorporate attribute names but exclude object names. This requires us to propose a query generation strategy that focuses on object attributes rather than exposing object identities (i.e., object classes). To achieve this, we perform query generation by hiding the object class of the object . We first replace the object classes with a substitution term . The substitution term , the relation , and the corresponding attribute are then concatenated to form the text query as:
3.4 Benchmark Statistics
@@
Table 2 shows the statistics of our OpenScan benchmark. We have collected eight linguistic aspects of attributes, providing a total of 153,644 attribute annotations across 341 attributes for 1,513 scenes in ScanNet200 [30]. In these aspects, the visual aspect of material is annotated manually, while other attributes are automatically generated via knowledge graph. There are 101.55 attribute annotations per scene in 1,513 3D indoor scenes. Besides, each object is annotated with an average of 3.15 attributes, indicating that most objects in ScanNet200 receive multiple attribute labels and are comprehensively represented. While certain linguistic aspects such as manner and synonym encompass a limited number of attributes, others like affordance and type consist of a wide range of attributes. We follow the training and validation split settings of ScanNet200.
@@
| OpenMask3D [34] | - | Mask3D [31] | SAM [18] |
| SAI3D [41] | - | - | Semantic-SAM [19] |
| MaskClustering [36] | - | - | CropFormer [26] |
| Open3DIS [24] | - | ISBNet [23] | Grounded-SAM [29] |
| OpenScene[25] | - | - | - |
| PLA[10] | ScanNet [8] | - | - |
| RegionPLC [37] | ScanNet [8] | - | - |
4 Experiments
@@
| OpenMask3D [34] | 7.2 | 7.5 | 8.5 | 12.8 | 16.9 | 13.0 | 12.2 | 18.8 | 9.9 | 15.4 |
| SAI3D [41] | 5.3 | 5.8 | 7.8 | 11.3 | 10.0 | 10.0 | 8.7 | 11.3 | 7.7 | 12.7 |
| MaskClustering [36] | 6.2 | 7.0 | 7.1 | 11.1 | 16.2 | 11.3 | 7.4 | 12.1 | 8.1 | 12.0 |
| Open3DIS [24] | 11.9 | 12.8 | 14.2 | 19.2 | 26.7 | 19.2 | 18.7 | 28.3 | 15.8 | 23.7 |
@@
| OpenMask3D [34] | 9.1 | 10.0 | 11.2 | 15.4 | 19.7 | 16.0 | 15.4 | 22.1 | 12.5 | 19.9 |
| SAI3D [41] | 8.4 | 8.3 | 11.4 | 15.7 | 16.7 | 15.3 | 13.6 | 17.1 | 11.6 | 18.8 |
| MaskClustering [36] | 10.7 | 12.3 | 13.3 | 18.4 | 30.3 | 21.8 | 13.5 | 20.6 | 14.6 | 23.3 |
| Open3DIS [24] | 14.8 | 16.0 | 17.9 | 22.3 | 30.6 | 24.1 | 21.9 | 33.6 | 19.3 | 29.4 |
@@
| OpenMask3D [34] | 10.4 | 11.6 | 13.0 | 17.4 | 20.6 | 18.9 | 17.1 | 25.0 | 14.2 | 23.1 |
| SAI3D [41] | 10.5 | 10.7 | 13.4 | 18.2 | 20.0 | 18.7 | 16.0 | 22.9 | 14.1 | 24.1 |
| MaskClustering [36] | 13.7 | 15.8 | 17.7 | 23.1 | 36.6 | 28.2 | 17.2 | 25.6 | 18.7 | 30.1 |
| Open3DIS [24] | 16.7 | 16.8 | 20.2 | 24.2 | 33.1 | 25.5 | 24.7 | 36.7 | 21.4 | 32.8 |
We conduct our experiments on the validation set of our OpenScan across eight linguistic aspects using the publicly available OV-3D models. For 3D instance segmentation, we evaluate OpenMask3D [34], SAI3D [41], MaskClustering [36], and Open3DIS [24]. For 3D semantic segmentation, we evaluate OpenScene [25], PLA [10], and RegionPLC [37]. Table 3 summarizes the information of these models: the training set, pre-trained 3D proposal, and pre-trained 2D proposal. All experiments are conducted on one RTX4090 GPU.
4.1 Main Results
The results of 3D instance segmentation on our OpenScan benchmark are presented in Table 4. We evaluate OpenMask3D [34], SAI3D [41], MaskClustering [36], and Open3DIS [24] across 341 attributes from our OpenScan and 198 object classes from ScanNet200 [30]. Results show that the performance of these OV-3D models on our GOV-3D benchmark, OpenScan, are significantly lower than those on the classic OV-3D dataset, ScanNet200. This gap underscores that our proposed GOV-3D task is a more challenging extension of the traditional OV-3D task.
When comparing the results of each OV-3D model across different linguistic aspects, we observe higher performances in the synonym and material aspects but struggle in the affordance and property aspects. The high performance in the synonym aspect can be attributed to the close similarity between attributes in this aspect and object classes, making recognition easier compared to the more abstract affordance and property aspects. An example of these closely related terms is shown in Figure 2, where the corresponding synonym aspect of the object class “nightstand” is “bedside table”. The high performance in the material aspect highlights the ability of these OV-3D models to recognize visual patterns. By utilizing CLIP [27] for 3D scene understanding, these models benefit from its visual patterns, including material and color from pre-trained image-text pairs, enhancing their comprehension of visual attributes beyond other attributes.
@@ 4.1 Main Results
| OpenScene [25] | 0.64 | 3.46 | 47.5 | 70.7 |
| PLA [10] | 0.03 | 4.25 | 66.6 | 77.5 |
| RegionPLC [37] | 0.25 | 4.23 | 68.7 | 78.7 |
@@
For 3D semantic segmentation, we evaluate OpenScene [25], PLA [10], and RegionPLC [37], reporting the average score of all attributes in our OpenScan and that of all object classes in ScanNet [8]. Table 5 shows that although these OV-3D models perform well in recognizing object classes, they exhibit poor performances on linguistic aspects with low mIoU and mAcc metrics. The methods for semantic segmentation suffer from a more significant performance drop on OpenScan when compared with those for instance segmentation. This drop can be caused by several factors. First, there is a significant discrepancy in the vocabulary size between ScanNet and our OpenScan. A larger vocabulary size implies a more diverse set of semantic concepts that the model needs to comprehend, making our OpenScan more challenging and practical in real-world scenarios. In addition, the arbitrary nature of object attributes in contrast to object classes adds complexity to the GOV-3D task. Besides, the lack of both robust 3D proposals (e.g., Mask3D [31]) and 2D proposals (e.g., SAM [18]) for class-agnostic masks can also be attributed to the drop. Conversely, instance segmentation models like OpenMask3D [34] leverage strong instance-level knowledge, e.g., proposals extracted from Mask3D and SAM, to effectively segment novel 3D objects, leading to higher performances on the GOV-3D task.
4.2 The Impact of Pre-trained Vocabulary Size

### v3

@@ Abstract
Introduction
Open-vocabulary 3D scene understanding (OV-3D) involves recognizing object classes that are not included in the training set. It is important to applications like autonomous driving (Bojarski et al. 2016) and robotics (Zeng et al. 2018). Recently, vision-language models (VLMs), e.g., CLIP (Radford et al. 2021), have achieved significant progress by leveraging large-scale image-text datasets with semantically rich captions. The impressive capability of VLMs in capturing rich contexts between images and texts has inspired further exploration in open-vocabulary tasks in both 2D (Gu et al. 2021; Zhong et al. 2022) and 3D (Takmaz et al. 2023; Peng et al. 2023) domains.
For AI systems, the capability to comprehend diverse linguistic aspects of object-related attributes and their association with corresponding objects is as important as the identification of the objects themselves. Consequently, the field of open-vocabulary 3D scene understanding should ideally extend beyond specific object classes to encompass complex object-related attributes, such as affordance and material, articulated through natural languages. However, the generalization ability of existing OV-3D methods (Peng et al. 2023; Takmaz et al. 2023; Yan et al. 2024; Yin et al. 2024; Nguyen et al. 2024) to object-related attributes has not been thoroughly and systematically explored. Besides, evaluating the ability of an OV-3D model to recognize specific object attributes is difficult due to the shortage of large-scale OV-3D attribute benchmarks. Existing OV-3D benchmarks, such as ScanNet (Dai et al. 2017) and ScanNet200 (Rozenberszki et al. 2022), primarily focus on object classes, and do not include annotations of object attributes for evaluating the generalized ability of OV-3D methods.
@@ Introduction
Existing 3D scene understanding benchmarks, such as ScanNet (Dai et al. 2017), ScanNet200 (Rozenberszki et al. 2022), and ScanNet++ (Yeshwanth et al. 2023), only provide annotations for object classes. To address this limitation of existing benchmarks, we construct a new benchmark, named OpenScan, for the GOV-3D task. OpenScan is constructed based on the ScanNet200 (Rozenberszki et al. 2022) benchmark. It expands the single category of object classes in ScanNet200 into eight linguistic aspects of object-related attributes, including affordance, property, type, manner, synonym, requirement, element, and material. This allows each object to be associated with some generalized knowledge beyond object classes. With our OpenScan benchmark, it becomes possible to comprehensively evaluate existing OV-3D models from various aspects, enabling a quantitative assessment of their generalization capabilities in understanding abstract object attributes.
@@
Related Work
@@
Open-Vocabulary 3D Scene Understanding. The study of open-vocabulary 3D scene understanding has been relatively limited compared to open-vocabulary 2D understanding. This is primarily due to the complexity and difficulty in obtaining 3D datasets. OpenMask3D (Takmaz et al. 2023) introduces the open-vocabulary 3D instance segmentation task. It proposes the first approach for the task in a zero-shot setting. OpenScene (Peng et al. 2023) also proposes a zero-shot method for open-vocabulary 3D scene understanding. Beyond the object class, it is able to utilize arbitrary text queries for semantic segmentation. Previous methods have mainly focused on object context for 3D scene understanding. PLA (Ding et al. 2023) and RegionPLC (Yang et al. 2024) extend the context to a more coarse-to-fine semantic representation to provide a more comprehensive supervision. Recently, OpenIns3D (Huang et al. 2024), Open3DIS (Nguyen et al. 2024), and SAI3D (Yin et al. 2024) utilize powerful 2D segmentation models to generate 2D instances and then merge them into 3D instances. Instead of utilizing accurate 2D masks from 2D segmentation models, MaskClustering (Yan et al. 2024) leverages clustering algorithms to perform zero-shot 3D segmentation. Recently, UniSeg3D (Xu et al. 2024) proposes a unified framework for 3D scene understanding. However, these methods only provide qualitative results for object attributes and lack a thorough evaluation of performance beyond object classes. This motivates us to conduct a quantitative evaluation that encompasses a wider range of object attributes.
@@
3D Scene Understanding Benchmark. Existing open-vocabulary 3D scene understanding benchmarks, e.g., ScanNet (Dai et al. 2017), ScanNet200 (Rozenberszki et al. 2022), S3DIS (Armeni et al. 2016), and Matterport3D (Chang et al. 2017), utilize RGB-D cameras, while ARKitScenes (Baruch et al. 2021), Replica (Straub et al. 2019), and ScanNet++ (Yeshwanth et al. 2023) leverage high-resolution laser scanners to capture high-fidelity 3D data for 3D reconstructions. Our proposed OpenScan benchmark expands the object class annotations of the open-vocabulary 3D scene understanding benchmarks (i.e., ScanNet200) to object-related attributes. Similar to the 3D referring (Chen, Chang, and Nießner 2020) and 3D reasoning (Huang et al. 2025) benchmarks, our OpenScan introduces new annotations for existing 3D scans to locate 3D objects via text queries. Unlike these tasks, which assume the queried object exists in the scene, our introduced GOV-3D task requires discriminative capabilities to determine whether the query presents in the scene. MMScan (Lyu et al. 2024) provides a benchmark for visual attribute understanding but lacks commonsense-related attribute annotations (e.g., “synonym” and “requirement”) included in our OpenScan. Recently, SceneFun3D (Delitzas et al. 2024) provides a large-scale 3D dataset with annotations for functionality and affordance interactions in 3D scenes. However, our OpenScan differs from SceneFun3D by considering a broader range of attributes. Specifically, while SceneFun3D focuses on function or affordance understanding, our OpenScan covers eight linguistic aspects, with affordance representing just one of them. Besides, while SceneFun3D focuses on element-level human-scene interaction (e.g., “door handle”), our OpenScan is designed for object-level scene understanding (e.g., “door”).
Task Setting and Benchmark
Task Formulation
@@
OV-3D. Let represent 3D scenes with points, denote RGB image frames, and is a vocabulary set of text sentences, each describing the object class that we aim to detect. An OV-3D model, , generates predictions with high confidence scores, . Predictions are then compared with the GT label for evaluation.
@@
GOV-3D. The existing 3D scene understanding benchmark, denoted as , comprises a collection of object-label pairs. Each pair consists of an object represented as a 3D mask and its corresponding class label . The benchmark is composed of multiple 3D scenes with points, and RGB image frames . Building upon this, GOV-3D extends the class label to object attribute . Accordingly, the attribute set for 3D scenes is defined as , composed of text sentences, each describing a specific attribute that we aim to detect. A GOV-3D model, , produces predictions with high confidence scores, . The evaluation of the GOV-3D task involves comparing the predictions with the GT label .
Benchmark Description
@@
The OpenScan benchmark is constructed based on the ScanNet200 (Rozenberszki et al. 2022) benchmark, which consists of 200 object classes with more than 1,500 3D scans. Since the ScanNet200 benchmark contains only an object-level class annotation for each object, it is not suitable for our GOV-3D task. To perform the GOV-3D task, we construct the OpenScan benchmark by leveraging the object annotations of the ScanNet benchmark. Our OpenScan provides attribute annotations for each object, expanding the single category of object classes in ScanNet200 into eight linguistic aspects of object-related attributes, including affordance, property, type, manner, synonym, requirement, element, and material. Figure 2 shows an example from our OpenScan benchmark. The target objects in our OpenScan are annotated with eight linguistic aspects of object attributes. The explanations of these eight object attributes are as follows:
@@
Benchmark Annotation
Object-Attribute Association with Knowledge Graph. We associate each object with various attributes using knowledge graphs, as illustrated in Figure 3. Let denotes the existing 3D scene understanding benchmark, e.g., ScanNet200 (Rozenberszki et al. 2022) in our implementation, where is a target object, is the corresponding class label, and denotes the number of target object and label pairs. The benchmark is composed of multiple 3D scenes with points. Let denotes the knowledge graph, where is the node set and is the edge set. The nodes are natural language words and phrases, and the edges are relation knowledge connecting them. Each edge is directional, and can be represented as a tuple , where are the names of the head node and the tail node, is the relation, and is the importance weight of this relation. We extract the relation knowledge from the popular and high-quality NLP knowledge base ConceptNet (Speer, Chin, and Havasi 2017). An example of relation knowledge from it is:
We query a set of relation knowledge linked to object class from the knowledge graph . Formally, for each edge within it, the head node name is the same as the input object class, i.e., . The query process is defined as:
These object-attribute pairs form the basic annotations of OpenScan, which is useful in the GOV-3D task. Finally, each 3D object is assigned a relation knowledge through annotations, serving as commonsense knowledge as:
Manual Annotation. For the visual attribute that cannot be inferred without human perception, we manually annotate each 3D object following a rigorous protocol. We create a web interface for annotators to select each object’s visual attribute. Specifically, for each scene, annotators view the 3D point cloud and the target’s corresponding 2D image. Taking the material attribute as an example, annotators are tasked with identifying the primary material composition of the target object. Any 3D object with an ambiguous appearance is carefully identified through different camera views of the scene and the corresponding image frames around the object. Finally, each 3D object is assigned with a relation and a visual attribute like material through annotations, serving as visual appearance in:
Attribute Classification. To better organize our benchmark, we manually group each attribute into eight linguistic aspects based on relation and attribute . This process involves considering the nature of the relation and attribute , and how they align with each linguistic aspect. Subsequently, each attribute is assigned to a linguistic aspect.
Attribute Verification. After the initial attribute classification, we manually verify each 3D object with its corresponding attribute and linguistic aspect. If a 3D object contains multiple attributes within a single linguistic aspect, we manually assign the most representative attribute to ensure evaluation consistency. This attribute is uniquely tied to the object class of to eliminate cross-class ambiguity. We also filter out similar attributes (e.g., “store things” and “store somethings”), preserving only one attribute to ensure consistency. During the verification process, the attributes across eight linguistic aspects are reduced from 528 to 341.
Query Generation. A practical GOV-3D query should incorporate attribute names but exclude object names, requiring a query strategy that focuses on attributes rather than exposing object identities (i.e., object classes). To achieve this, we perform query generation by hiding the object class of the object . We first replace the object classes with a substitution term . The substitution term , the relation , and the corresponding attribute are then concatenated to form the text query as:
@@
Benchmark Statistics
@@
Table 1 shows the statistics of our OpenScan benchmark. We have collected eight linguistic aspects of attributes, providing a total of 153,644 attribute annotations across 341 attributes for 1,513 scenes in ScanNet200 (Rozenberszki et al. 2022). In these aspects, the visual aspect of material is annotated manually, while other attributes are automatically generated via knowledge graph. There are 101.55 attribute annotations per scene in 1,513 3D indoor scenes. Besides, each object is annotated with an average of 3.15 attributes, indicating that most objects in ScanNet200 receive multiple attribute labels and are comprehensively represented. While certain linguistic aspects such as manner and synonym encompass a limited number of attributes, others like affordance and type consist of a wide range of attributes. We follow the training and validation split settings of ScanNet200.
Evaluation Metrics
We employ commonly used OV-3D metrics to evaluate our GOV-3D task. For semantic segmentation, we follow (Peng et al. 2023; Ding et al. 2023; Yang et al. 2024) to apply mean IoU (mIoU) and mean accuracy (mAcc). For instance segmentation, we follow (Takmaz et al. 2023; Yin et al. 2024; Yan et al. 2024; Nguyen et al. 2024) to apply average precision (AP) at IoU scores of 25% (AP25), 50% (AP50), and the mean of AP from 50% to 95% at 5% steps. We compute the mean score (Mean) across all attributes to obtain the overall performance.
Experiments
@@
| OpenMask3D | - | Mask3D | SAM |
| SAI3D | - | - | Semantic-SAM |
| MaskClustering | - | - | CropFormer |
| Open3DIS | - | ISBNet | Grounded-SAM |
| OpenScene | - | - | - |
| PLA | ScanNet | - | - |
| RegionPLC | ScanNet | - | - |
We conduct our experiments on the validation set of our OpenScan across eight linguistic aspects using the publicly available OV-3D models. All OV-3D models are evaluated in a zero-shot setting without training on the OpenScan benchmark. For 3D instance segmentation, we evaluate OpenMask3D (Takmaz et al. 2023), SAI3D (Yin et al. 2024), MaskClustering (Yan et al. 2024), and Open3DIS (Nguyen et al. 2024). For 3D semantic segmentation, we evaluate OpenScene (Peng et al. 2023), PLA (Ding et al. 2023), and RegionPLC (Yang et al. 2024). Table 2 summarizes the information of these models: the training set, pre-trained 3D proposal, and pre-trained 2D proposal. All experiments are conducted on one NVIDIA RTX 4090 GPU.
Main Results
3D Instance Segmentation. We evaluate OpenMask3D (Takmaz et al. 2023), SAI3D (Yin et al. 2024), MaskClustering (Yan et al. 2024), and Open3DIS (Nguyen et al. 2024) across 341 attributes from our OpenScan and 198 object classes from ScanNet200 (Rozenberszki et al. 2022). Table 3 show that the performance of these OV-3D models on our GOV-3D benchmark, OpenScan, are significantly lower than those on the classic OV-3D dataset, ScanNet200. This gap underscores that our proposed GOV-3D task is a more challenging extension of the traditional OV-3D task.
When comparing the results of each OV-3D model across different linguistic aspects, we observe higher performances in the synonym and material aspects but struggle in the affordance and property aspects. The high performance in the synonym aspect can be attributed to the close similarity between attributes in this aspect and object classes, making recognition easier compared to the more abstract affordance and property aspects. An example of these closely related terms is shown in Figure 2, where the corresponding synonym aspect of the object class “nightstand” is “bedside table”. The high performance in the material aspect highlights the ability of these OV-3D models to recognize visual patterns. By utilizing CLIP (Radford et al. 2021) for 3D scene understanding, these models benefit from its visual patterns, including material and color from image-text pretraining , enhancing their comprehension of visual attributes beyond other attributes.
@@
| OpenMask3D (Takmaz et al. 2023) | 7.2 | 7.5 | 8.5 | 12.8 | 16.9 | 13.0 | 12.2 | 18.8 | 9.9 | 15.4 |
| SAI3D (Yin et al. 2024) | 5.3 | 5.8 | 7.8 | 11.3 | 10.0 | 10.0 | 8.7 | 11.3 | 7.7 | 12.7 |
| MaskClustering (Yan et al. 2024) | 6.2 | 7.0 | 7.1 | 11.1 | 16.2 | 11.3 | 7.4 | 12.1 | 8.1 | 12.0 |
| Open3DIS (Nguyen et al. 2024) | 11.9 | 12.8 | 14.2 | 19.2 | 26.7 | 19.2 | 18.7 | 28.3 | 15.8 | 23.7 |
@@
| OpenMask3D (Takmaz et al. 2023) | 9.1 | 10.0 | 11.2 | 15.4 | 19.7 | 16.0 | 15.4 | 22.1 | 12.5 | 19.9 |
| SAI3D (Yin et al. 2024) | 8.4 | 8.3 | 11.4 | 15.7 | 16.7 | 15.3 | 13.6 | 17.1 | 11.6 | 18.8 |
| MaskClustering (Yan et al. 2024) | 10.7 | 12.3 | 13.3 | 18.4 | 30.3 | 21.8 | 13.5 | 20.6 | 14.6 | 23.3 |
| Open3DIS (Nguyen et al. 2024) | 14.8 | 16.0 | 17.9 | 22.3 | 30.6 | 24.1 | 21.9 | 33.6 | 19.3 | 29.4 |
@@
| OpenMask3D (Takmaz et al. 2023) | 10.4 | 11.6 | 13.0 | 17.4 | 20.6 | 18.9 | 17.1 | 25.0 | 14.2 | 23.1 |
| SAI3D (Yin et al. 2024) | 10.5 | 10.7 | 13.4 | 18.2 | 20.0 | 18.7 | 16.0 | 22.9 | 14.1 | 24.1 |
| MaskClustering (Yan et al. 2024) | 13.7 | 15.8 | 17.7 | 23.1 | 36.6 | 28.2 | 17.2 | 25.6 | 18.7 | 30.1 |
| Open3DIS (Nguyen et al. 2024) | 16.7 | 16.8 | 20.2 | 24.2 | 33.1 | 25.5 | 24.7 | 36.7 | 21.4 | 32.8 |
@@
| OpenScene (Peng et al. 2023) | 0.45 | 1.87 | 47.5 | 70.7 |
| PLA (Ding et al. 2023) | 0.01 | 2.37 | 66.6 | 77.5 |
| RegionPLC (Yang et al. 2024) | 0.07 | 2.36 | 68.7 | 78.7 |
@@
3D Semantic Segmentation. We evaluate OpenScene (Peng et al. 2023), PLA (Ding et al. 2023), and RegionPLC (Yang et al. 2024), reporting the average score of all attributes in our OpenScan and that of all object classes in ScanNet (Dai et al. 2017). Table 4 shows that although these OV-3D models perform well in recognizing object classes, they exhibit poor performances on linguistic aspects with low mIoU and mAcc metrics. The methods for semantic segmentation suffer from a more significant performance drop on OpenScan when compared with those for instance segmentation. This drop can be caused by several factors. First, there is a significant discrepancy in the vocabulary size between ScanNet and our OpenScan. A larger vocabulary size implies a more diverse set of semantic concepts that the model needs to comprehend, making our OpenScan more challenging and practical in real-world scenarios. In addition, the arbitrary nature of object attributes in contrast to object classes adds complexity to the GOV-3D task. Besides, the lack of both robust 3D proposals (e.g., Mask3D (Schult et al. 2023)) and 2D proposals (e.g., SAM (Kirillov et al. 2023)) for class-agnostic masks can also be attributed to the drop. Conversely, instance segmentation models like OpenMask3D (Takmaz et al. 2023) leverage strong instance-level knowledge, e.g., proposals extracted from Mask3D and SAM, to effectively segment novel 3D objects, leading to higher performances on the GOV-3D task.
| Method | Template | AP | AP50 | AP25 |
| OpenMask3D (Takmaz et al. 2023) | - | 9.7 | 12.2 | 14.1 |
| ✓ | 9.9 | 12.5 | 14.2 | |
| SAI3D (Yin et al. 2024) | - | 6.7 | 10.1 | 12.8 |
| ✓ | 7.7 | 11.6 | 14.1 | |
| MaskClustering (Yan et al. 2024) | - | 6.8 | 12.0 | 14.6 |
| ✓ | 8.1 | 14.6 | 18.7 | |
| Open3DIS (Nguyen et al. 2024) | - | 15.6 | 19.2 | 21.3 |
| ✓ | 15.8 | 19.3 | 21.4 |
The Impact of the Pre-trained Vocabulary Size